OpenStudy (anonymous):

y=x^2+3x-10 Find the x intercepts of the graph?

OpenStudy (anonymous):

x intercepts are when y is equal to 0 so 0 - x^2 + 3x -10 (x+5)(x-2) x = -5, 2

OpenStudy (anonymous):

* 0 = x^2 + 3x -10

hero (hero):

I have no idea how you got your result manda. All you had to do was set y = 0 and solve for x.

OpenStudy (anonymous):

i did. you did -5x +2x but that doesn't add up to 3, it adds up to -3 and the equation says +3x

OpenStudy (anonymous):

check your work hero

hero (hero):

x^2+5x-2x -10=0 x(x+5)-2(x+5) =0 (x+5)(x-2) =0 x = 2,-5

hero (hero):

Somehow, I convinced myself that the way I did it the first time was correct